OpenStudy (anonymous):

Approximately 15% of peole are left-handed. If two people are selected at random, what is the probability of the following events? P(At least one is right-handed)? Is this correct? 1-P(none are RH) so..1-P(LH and LH) 1-(.85)(.85) =1-.7225 =.2775...is this correct?

OpenStudy (amistre64):

should we assume that people can either be left or right handed, but not both?

OpenStudy (anonymous):

well I guess it's possible to be both

OpenStudy (anonymous):

but I've never had a question ask me that...

OpenStudy (amistre64):

since no stats are given as to ambidextrous peoples ... P(l) = .15, P(r) = .85 at least one is P(l) lr + ll

OpenStudy (amistre64):

at least one is right handed that is .... lr + rr

OpenStudy (anonymous):

Thank you! that makes sense.

OpenStudy (amistre64):

the possible ways to choose 2 people are: ll .15*.15 lr .15*.85 rl .85*.15 rr .85*.85 at least 1 r in it is: lr + rl + rr does that sound right?

OpenStudy (amistre64):

1 - ll :) .9775

OpenStudy (anonymous):

yes it does. makes so much more sense. Thank you!

OpenStudy (anonymous):

wait so is the answer .85?

OpenStudy (amistre64):

ll+lr+rl+rr = 1 lr+rl+rr = 1 - ll since l = .15 ... 1 - .15^2 = .9775

OpenStudy (anonymous):

oh sorry I forgot the one part..thank you!